# Dual Frequency Branch Framework with Reconstructed Sliding Windows Attention for AI-Generated Image Detection

### v2

@@ Abstract
The rapid advancement of Generative Adversarial Networks (GANs) and diffusion models has enabled the creation of highly realistic synthetic images, presenting significant societal risks, such as misinformation and deception. As a result, detecting AI-generated images has emerged as a critical challenge. Existing researches emphasize extracting fine-grained features to enhance detector generalization, yet they often lack consideration for the importance and interdependencies of internal elements within local regions and are limited to a single frequency domain, hindering the capture of general forgery traces. To overcome the aforementioned limitations, we first utilize a sliding window to restrict the attention mechanism to a local window, and reconstruct the features within the window to model the relationships between neighboring internal elements within the local region. Then, we design a dual frequency domain branch framework consisting of four frequency domain subbands of DWT and the phase part of FFT to enrich the extraction of local forgery features from different perspectives. Through feature enrichment of dual frequency domain branches and fine-grained feature extraction of reconstruction sliding window attention, our method achieves superior generalization detection capabilities on both GAN and diffusion model-based generative images. Evaluated on diverse datasets comprising images from 65 distinct generative models, our approach achieves a 2.13% improvement in detection accuracy over state-of-the-art methods.
Index Terms:
Deepfake detection, AI-generated image detection, AI security, Window attention.I Introduction
@@
With the rapid advancement of deep learning, generative models such as GANs and diffusion models have revolutionized content creation, finding widespread applications across industries including entertainment and media. However, this progress has also introduced significant challenges. The increasing sophistication of deepfake technology has allowed the generation of highly realistic fake images and videos, making them increasingly difficult to detect and posing serious risks to societal trust, individual privacy, and security.
@@
To prevent the abuse of generative methods, a series of AI-generated image detection mechanisms have been conceived. Early methods primarily focused on analyzing image features, such as texture, color, and lighting [1, 2] to identify forgery traces. However, these methods are usually limited to the domain consistent scenario, where the training and testing sets come from the same type of generative method, resulting in poor generalized detection performance in cross-generative method scenarios. With the continuous deepening of research, it is widely recognized that artifacts are often confined to subtle and localized areas. Inspired by this, extensive research has been conducted to extract generalizable and fine-grained forgery features. Specifically, Tan et al. [3] extracted more generalized artifact information from the generator’s perspective by analyzing the relationship between adjacent pixels. Other works [4, 5, 6] have further amplified the local artifacts by means of segmenting patches, thus achieving better detection performance. In addition, some studies [7, 8, 9] captured forgery traces from the perspective of frequency domain to obtain more generalized detection performance.
@@
Therefore, we design a dual frequency branch framework consisting of four frequency bands in the DWT domain and the phase part in the FFT domain to extract richer local artifact features from multiple perspectives. Based on this, a reconstructed sliding window attention mechanism is designed to obtain more fine-grained and generalized local artifact features, thus improving the detection performance across various generation methods. To fully assess the generalizability of our method, we evaluate it on large image datasets generated by 65 distinct generation models111AttGAN, BEGAN, CramerGAN, InfoMaxGAN, MMDGAN, RelGAN, S3GAN, SNGAN, STGAN, ADM, DDPM, IDDPM, LDM, PNDM, VQDiffusion, SDv1, SDv2, Guided, Glide_50_27, Glide_100_10, Glide_100_27, LDM_100, LDM_200, LDM_200_cfg, DALLE, ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, ADM, Glide, Midjourney, SDv1.4, SDv1.5, VQDM, Wukong, DALLE2, R3GAN, StyleGAN3, StyleGAN-XL, StyleSwim, BlendFace, E4S, FaceSwap, InSwap, SimSwap, FLUX1-dev, Midjourney V6, GLIDE, DALLE-3, Imagen3, SD3, SDXL, BLIP, Infinite-ID, InstantID, IP-adapter, PhotoMaker, SocialRF, CommunityAI. Comprehensive experimental results show that our method achieves a significant improvement over state-of-the-art methods C2P_CLIP [17] and AIDE [18] by 5.8% and 7.0%. The overall contributions of the proposed method can be summarized as follows.
@@ II Related Work
With the continuous development of generation methods, including GAN[19, 20, 21, 22], diffusion [23, 24, 25, 26], customized generation [27, 28, 29, 30, 31] and so on, the generated images are becoming more and more realistic, and the naked eye can no longer distinguish the authenticity of the images. In order to solve the future brought about by the above problems, many researchers are committed to the detection of AI-generated images. In this section, we will briefly introduce the existing detection methods from two aspects.
II-A Exploration of Fine-Grained Features in AI-Generated Image Detection
@@
Early researches on AI-generated image detection often focused on specific facial regions, such as the eyes and lips, to identify forgery traces [32, 33]. However, as forgery techniques have advanced, these simple traces have been increasingly refined and obscured, rendering biometric-based detection methods insufficient for identifying sophisticated forgery images. Experiments show that forgery traces are often embedded in fine details, which motivates people to turn to extracting local artifact information, including spatial and frequency domains. In terms of spatial domains, some researches narrow the focus to local areas to extract artifact information, including limited receptive fields [34, 35], cropping the image into smaller patches [4, 36, 6, 5], etc. Specifically, Chai et al. [34] employ limited receptive fields to identify patches that render images detectable, Zheng et al. [6] divides the image into multiple patches and shuffles them, and trains a block-based convolutional network for feature extraction. In addition, LGrad [8] employs pretrained CNN models as transform functions to convert images into gradients and leverages these gradients to present universal artifacts. NPR [3] improve generalization by identifying upsampling artifacts, such as pixel grid distortions, in nearby regions. In the frequency domain, F3-Net [37] introduces frequency component division and the frequency statistical difference between real images and forgery images into face forgery detection. FreDect [38] observes significant artifacts in the frequency domain of GAN-generated images, attributed to the upsampling operation in GAN architectures. Luo et al. [39] utilize multiple high-frequency features of images to enhance generalization performance. FreqNet [9] utilize the Fast Fourier Transform (FFT) to extract global high-frequency information, revealing refined forgery features such as periodic noise or anomalies invisible in the spatial domain. AIDE [18] improved the generalization ability of forgery detection by extracting the highest and lowest fine-grained patches through DCT to capture low-level fine-grained artifact information.
II-B Application of Attention in AI-generated Image Detection
@@
Previous studies [10, 11, 12] has proven that: The attention mechanism can not only be trained to focus on important areas of the image, but also capture the long-range dependencies of the image, which is very useful in AI-generated image detection. Many works have explored this: Zhao et al.[40] employ multiple spatial attention heads, allowing the model to attend to distinct regions—such as the eyes and mouth—simultaneously, enhancing its ability to spot inconsistencies. Wu et al.[41] introduce a local attention module interacting with distant structures (e.g., PPG graphs), though further details from the original study would clarify this mechanism. In recent years, the attention mechanism has been widely integrated into advanced models such as Vision Transformer (ViT) [12] and CLIP [42], which not only improves the accuracy of detection, but also brings new possibilities for dealing with increasingly complex forgery techniques. For example, Ojha et al. [43] directly utilizes image features from the CLIP model for linear classification, LASTED [44] proposes designing textual labels to supervise the CLIP vision model through image-text contrastive learning, further advancing the field of AIGC detection, Fatformer [15] introduces a forgery-aware adapter to discern and integrate local forgery traces based on CLIP. CLIPMoLE [45] adapts a combination of shared and separate LoRAs within an MoE-based structure in deeper ViT blocks. What’s more, C2P_CLIP [17] explored the reasons and potential of CLIP’s effectiveness in AI-generated image detection, and for the first time integrated category common cues into the text encoder to inject category-related concepts into the image encoder, thereby improving detection performance.
@@
Existing attention mechanisms, such as Vision Transformers (ViT) and CLIP, typically capture global weights and long-range dependencies to extract broad semantic information. However, they lack sensitivity to the significance and interrelationships of elements within local regions, hindering their ability to effectively detect fine-grained forgery traces. Therefore, we propose a reconstructed sliding window attention mechanism, which enhancing the feature extraction capacity of the window attention block to model the importance and dependencies of local forgery features. In addition, a dual frequency branch network is designed to enrich the features within the window from different angles, which promotes the extraction of more comprehensive local forgery features in the window.
@@ III-B Reconstructed Sliding Window Attention Block
The Reconstructed Sliding Window Attention Block (RSWAttention) consists of two main components: (i) the Local Attention Block (LoAttention) and (ii) the Global Window MLP Layer (GMLP). Before feeding the features into LoAttention and GMLP, we preprocess them to enhance their representational quality. A convolution layer is applied to extract cross-channel features, following by a depthwise separable convolution to better capture the local features within each channel. The resulting convolved features are denoted as .
@@
III-C Dual Frequency Branch Feature Enrichment Module
where represent the low-frequency component of the input and the high-frequency components in the vertical, horizontal, and diagonal directions, respectively. To obtain more comprehensive information, we fuse the data from the four frequency domains, resulting in new input features . As illustrated in Figure 3, to extract detailed features, we divide each channel’s features into windows, with each window capturing localized spatial-frequency information within its respective frequency band. Each window is treated as a small square and tiled into rows within each band. This arrangement ensures that features within the windows are preserved in a structured format. Then we tile the features in the window here by rows. At this time, the features from each frequency band are tiled into one row in a specific manner, such that the four adjacent feature values correspond to the window from the original image. Additionally, as shown in Figure 3, we integrate features from all frequency bands to generate new input features . To further enhance the network’s performance, we apply LayerNorm [47], which ensures that the input data remains within a stable range, reduces dependency on initialization, and improves the stability of model training.
| CNNDetection | 54.9 | 66.6 | 54.2 | 76.0 | 53.3 | 72.9 | 53.0 | 71.3 | 51.9 | 63.7 | 52.0 | 64.5 | 51.6 | 63.1 | 51.8 | 61.3 | 52.8 | 67.4 |
| CLIPDetection | 69.7 | 87.6 | 79.1 | 99.8 | 77.9 | 94.5 | 78.5 | 95.3 | 95.0 | 99.3 | 99.0 | 99.3 | 74.0 | 92.5 | 87.3 | 97.5 | 82.6 | 95.7 |
| C2P_CLIP | 69.1 | 94.1 | 95.3 | 99.8 | 96.1 | 99.8 | 95.3 | 99.8 | 99.3 | 100.0 | 99.3 | 100.0 | 97.3 | 99.8 | 98.6 | 99.1 | 93.8 | 99.2 |
| CNNDetection | 97.9 | 99.9 | 66.1 | 93.5 | 67.0 | 97.1 | 55.9 | 80.5 | 70.1 | 97.0 | 76.5 | 91.5 | 72.9 | 93.2 | 51.7 | 66.5 | 62.1 | 89.4 |
| CLIPDetection | 100.0 | 100.0 | 99.7 | 100.0 | 75.7 | 97.8 | 95.1 | 99.3 | 84.4 | 97.1 | 98.7 | 99.8 | 99.9 | 99.4 | 67.4 | 82.0 | 64.2 | 68.7 |
| C2P_CLIP | 100.0 | 100.0 | 99.2 | 100.0 | 95.6 | 99.9 | 99.1 | 100.0 | 96.4 | 99.5 | 97.3 | 100.0 | 99.6 | 100.0 | 93.8 | 98.6 | 94.8 | 99.5 |
| CNNDetection | 52.1 | 67.8 | 51.7 | 67.1 | 51.7 | 64.7 | 50.6 | 59.2 | 50.6 | 59.2 | 51.5 | 63.5 | 50.5 | 58.5 | 50.1 | 49.2 | 61.3 | 76.7 |
| CLIPDetection | 67.6 | 87.2 | 61.7 | 85.5 | 56.7 | 74.5 | 63.1 | 87.2 | 62.8 | 86.3 | 84.9 | 96.8 | 71.2 | 90.9 | 49.5 | 60.2 | 73.5 | 83.6 |
| C2P_CLIP | 77.2 | 95.8 | 88.5 | 98.8 | 59.2 | 83.0 | 82.8 | 97.3 | 82.8 | 97.2 | 87.2 | 98.4 | 80.1 | 95.5 | 65.1 | 92.9 | 88.2 | 92.6 |
| CNNDetection | 97.8 | 100.0 | 49.9 | 48.4 | 64.0 | 91.6 | 49.7 | 60.0 | 58.0 | 88.9 | 62.1 | 89.4 | 49.3 | 34.2 | 49.6 | 35.2 | 50.0 | 44.3 |
| CLIPDetection | 99.5 | 100.0 | 93.0 | 98.6 | 68.4 | 79.0 | 87.1 | 95.0 | 94.5 | 99.3 | 91.0 | 97.8 | 47.2 | 42.7 | 72.4 | 82.5 | 69.3 | 79.4 |
| C2P_CLIP | 100.0 | 100.0 | 95.4 | 99.8 | 93.8 | 98.5 | 94.1 | 98.1 | 95.2 | 100.0 | 94.8 | 99.5 | 55.2 | 60.5 | 69.8 | 76.1 | 73.5 | 82.2 |
| CNNDetection | 50.1 | 44.3 | 49.7 | 45.3 | 50.3 | 50.3 | 49.9 | 39.3 | 51.5 | 67.3 | 50.4 | 53.8 | 51.0 | 50.3 | 52.8 | 66.3 | 52.6 | 65.5 |
| CLIPDetection | 48.9 | 48.2 | 50.1 | 51.5 | 46.3 | 39.6 | 44.8 | 35.6 | 69.4 | 81.7 | 46.8 | 38.4 | 48.1 | 44.1 | 49.0 | 48.6 | 50.7 | 53.6 |
| C2P_CLIP | 59.5 | 69.0 | 67.9 | 78.5 | 47.6 | 46.8 | 51.2 | 50.2 | 85.1 | 93.4 | 64.4 | 75.4 | 45.2 | 46.5 | 53.5 | 65.6 | 59.3 | 76.0 |
| CNNDetection | 49.8 | 44.7 | 50.8 | 55.8 | 50.1 | 67.1 | 50.0 | 50.1 | 49.6 | 51.5 | 50.3 | 46.3 | 49.9 | 50.5 | 53.6 | 57.6 |
| CLIPDetection | 57.8 | 68.9 | 72.4 | 84.2 | 67.4 | 79.5 | 62.3 | 73.9 | 47.0 | 41.4 | 49.2 | 46.2 | 49.9 | 50.0 | 63.3 | 66.4 |
| C2P_CLIP | 84.9 | 93.8 | 85.2 | 93.4 | 86.8 | 93.9 | 55.1 | 65.9 | 50.6 | 52.0 | 56.4 | 62.3 | 51.0 | 54.5 | 71.0 | 77.3 |

### v3

@@ Abstract
The rapid advancement of Generative Adversarial Networks (GANs) and diffusion models has enabled the creation of highly realistic synthetic images, presenting significant societal risks, such as misinformation and deception. As a result, detecting AI-generated images has emerged as a critical challenge. Existing research emphasizes extracting fine-grained features to enhance detector generalization, yet they often lack consideration for the importance and interdependencies of internal elements within local regions and are limited to a single frequency domain, hindering the capture of general forgery traces. To overcome the aforementioned limitations, we first utilize a sliding window to restrict the attention mechanism to a local window, and reconstruct the features within the window to model the relationships between neighboring internal elements within the local region. Then, we design a dual frequency domain branch framework consisting of four frequency domain subbands of DWT and the phase part of FFT to enrich the extraction of local forgery features from different perspectives. Through feature enrichment of dual frequency domain branches and fine-grained feature extraction of reconstruction sliding window attention, our method achieves superior generalization detection capabilities on both GAN and diffusion model-based generative images. Evaluated on diverse datasets comprising images from 65 distinct generative models, our approach achieves a 2.13% improvement in detection accuracy over state-of-the-art methods. Our code is available at https://github.com/HorizonTEL/DFFreq-main.
I Introduction
@@
With the rapid advancement of deep learning, generative models [25, 12, 33] such as GANs and diffusion models have revolutionized content creation, finding widespread applications across industries including entertainment and media. However, this progress has also introduced significant challenges [64, 59, 58]. The increasing sophistication of deepfake technology has allowed the generation of highly realistic fake images and videos, making them increasingly difficult to detect and posing serious risks to societal trust, individual privacy, and security [32, 34, 62, 63].
@@
To prevent the abuse of generative methods, a series of AI-generated image detection mechanisms have been conceived. Early methods primarily focused on analyzing image features, such as texture, color, and lighting [79, 39] to identify forgery traces. However, these methods are usually limited to the domain consistent scenario, where the training and testing sets come from the same type of generative method, resulting in poor generalized detection performance in cross-generative method scenarios. With the continuous deepening of research, it is widely recognized that artifacts are often confined to subtle and localized areas. Inspired by this, extensive research has been conducted to extract generalizable and fine-grained forgery features [57, 45, 55, 54, 76, 74, 6, 73, 17, 46, 8]. Specifically, Tan et al. [56] extracted more generalized artifact information from the generator’s perspective by analyzing the relationship between adjacent pixels. Other works [9, 2, 82] have further amplified the local artifacts by means of segmenting patches, thus achieving better detection performance. In addition, some studies [23, 57, 55] captured forgery traces from the perspective of frequency domain to obtain more generalized detection performance.
@@
Therefore, we design a dual frequency branch framework consisting of four frequency bands in the DWT domain and the phase part in the FFT domain to extract richer local artifact features from multiple perspectives. Based on this, a reconstructed sliding window attention mechanism is designed to obtain more fine-grained and generalized local artifact features, thus improving the detection performance across various generation methods. To fully assess the generalizability of our method, we evaluate it on large image datasets generated by 65 distinct generation models111AttGAN, BEGAN, CramerGAN, InfoMaxGAN, MMDGAN, RelGAN, S3GAN, SNGAN, STGAN, ADM, DDPM, IDDPM, LDM, PNDM, VQDiffusion, SDv1, SDv2, Guided, Glide_50_27, Glide_100_10, Glide_100_27, LDM_100, LDM_200, LDM_200_cfg, DALLE, ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, ADM, Glide, Midjourney, SDv1.4, SDv1.5, VQDM, Wukong, DALLE2, R3GAN, StyleGAN3, StyleGAN-XL, StyleSwim, BlendFace, E4S, FaceSwap, InSwap, SimSwap, FLUX1-dev, Midjourney V6, GLIDE, DALLE-3, Imagen3, SD3, SDXL, BLIP, Infinite-ID, InstantID, IP-adapter, PhotoMaker, SocialRF, CommunityAI. Comprehensive experimental results show that our method achieves a significant improvement over state-of-the-art methods C2P-CLIP [54] and AIDE [74] by 5.8% and 7.0%. The overall contributions of the proposed method can be summarized as follows.
@@ II Related Work
With the continuous development of generation methods, including GAN[24, 3, 11, 51], diffusion [12, 50, 60, 27], customized generation [65, 70, 31, 29, 78] and so on, the generated images are becoming increasingly realistic, and the naked eye can no longer distinguish the authenticity of the images. In order to solve the future issues mentioned above, many researchers are committed to detecting AI-generated images. In this section, we will briefly introduce the existing detection methods.
II-A Deepfake Detection
Early research on deepfake detection often focused on specific facial regions, such as the eyes and lips, to identify forgery traces [18, 7]. However, as forgery techniques have advanced, these simple traces have increasingly been refined and obscured, rendering biometric-based detection methods insufficient for identifying sophisticated forgery images. To enhance cross-domain performance, a variety of universal detectors have been proposed based on frequency analysis [47, 15, 41], wavelet [43], forgery augmentations [71, 53, 22], ID information [13, 20], and decoupled representation learning [36, 77, 75, 30]. Specifically, UCF [77] extracts common forgery features across manipulation techniques; LSDA [75] constructs and simulates variations within and across forgery features in the latent space, expanding the forgery feature space and enabling the learning of a more generalizable decision boundary. FPG [71] is proposed to elevate the generalization of deepfake detectors from the perspective of real-time perception analysis, while Li et al. propose a critical forgetting mechanism to force the detector to pay more attention to common subtle forgery traces that were previously overlooked, thereby enhancing the generalization capability.
II-B Exploration of Fine-Grained Features in AI-Generated Image Detection
@@
Experiments show that forgery traces are often embedded in fine details, which motivates people to extract local artifact information, including spatial and frequency domains. In terms of spatial domains, some research narrows the focus to local areas to extract artifact information, including limited receptive fields [5, 4], cropping the image into smaller patches [9, 10, 82, 2], etc. Specifically, Chai et al. [5] employ limited receptive fields to identify patches that render images detectable; Zheng et al. [82] divide the image into multiple patches, and train a block-based convolutional network for feature extraction. In addition, LGrad [57] employs pretrained CNN models as transform functions to convert images into gradients and leverages these gradients to present universal artifacts. NPR [56] improves generalization by identifying upsampling artifacts, such as pixel grid distortions, in nearby regions. In the frequency domain, F3-Net [47] introduces frequency component division and the frequency statistical difference between real images and forgery images into face forgery detection. FreDect [15] observes significant artifacts in the frequency domain of GAN-generated images, attributed to the upsampling operation in GAN architectures. Luo et al. [41] utilize multiple high-frequency features of images to enhance generalization performance. FreqNet [55] utilizes the Fast Fourier Transform to extract global high-frequency information, revealing refined forgery features such as periodic noise or anomalies invisible in the spatial domain. AIDE [74] improved the generalization ability of forgery detection by extracting the highest and lowest fine-grained patches through DCT to capture low-level, fine-grained artifact information. What’s more, some methods learn more generalizable artifact representations through large-scale data and augmentation strategies. For example, DRCT [6] leverages diffusion-based reconstruction to achieve better semantic alignment, while B-Free [17] mitigates dataset bias using self-conditioned inpainted reconstructions and content-aware augmentation.
II-C Application of Attention in AI-generated Image Detection
@@
Previous studies [72, 61, 14] has proven that: The attention mechanism can not only be trained to focus on important areas of the image, but also capture the long-range dependencies of the image, which is very useful in AI-generated image detection. Many works have explored this: Zhao et al.[81] employ multiple spatial attention heads, allowing the model to attend to distinct regions—such as the eyes and mouth—simultaneously, enhancing its ability to spot inconsistencies. Wu et al.[69] introduce a local attention module interacting with distant structures (e.g., PPG graphs), though further details from the original study would clarify this mechanism. In recent years, the attention mechanism has been widely integrated into advanced models such as Vision Transformer (ViT) [14] and CLIP [48], which not only improve the accuracy of detection, but also bring new possibilities for dealing with increasingly complex forgery techniques. For example, Ojha et al. [45] directly utilize image features from the CLIP model for linear classification; LASTED [68] proposes designing textual labels to supervise the CLIP vision model through image-text contrastive learning, further advancing the field of AIGC detection; Fatformer [37] introduces a forgery-aware adapter to discern and integrate local forgery traces based on CLIP. CLIPMoLE [40] adapts a combination of shared and separate LoRAs within an MoE-based structure in deeper ViT blocks. What’s more, C2P-CLIP [54] explored the reasons and potential of CLIP’s effectiveness in AI-generated image detection, and integrated category common cues into the text encoder to inject category-related concepts into the image encoder, thereby improving detection performance.
@@
Existing attention mechanisms, such as Vision Transformers and CLIP, typically capture global weights and long-range dependencies to extract broad semantic information. However, they lack sensitivity to the significance and interrelationships of elements within local regions, hindering their ability to effectively detect fine-grained forgery traces. Therefore, we propose a reconstructed sliding window attention mechanism that enhances the feature extraction capacity of the window attention block to model the importance and dependencies of local forgery features. In addition, a dual frequency branch network is designed to enrich the features within the window from different angles, which promotes the extraction of more comprehensive local forgery features in the window.
@@ III-B Reconstructed Sliding Window Attention Block
The Reconstructed Sliding Window Attention Block (RSWAttention) consists of two main components: (i) the Local Attention Block (LoAttention) and (ii) the Global Window MLP Layer (GMLP). Before feeding the features into LoAttention and GMLP, we preprocess them to enhance their representational quality. A convolution layer is applied to extract cross-channel features, followed by a depthwise separable convolution to better capture the local features within each channel. The resulting convolved features are denoted as .
@@
III-C Dual Frequency Branch Feature Enrichment Module
where represents the input image, represents the low-frequency component of the input and the high-frequency components in the vertical, horizontal, and diagonal directions, respectively. To obtain more comprehensive information, we fuse the data from the four frequency domains, resulting in new input features . As illustrated in Figure 3, to extract detailed features, we divide each channel’s features into windows, with each window capturing localized spatial-frequency information within its respective frequency band. Each window is treated as a small square and tiled into rows within each band. This arrangement ensures that features within the windows are preserved in a structured format. Then we tile the features in the window here by rows. At this time, the features from each frequency band are tiled into one row in a specific manner, such that the four adjacent feature values correspond to the window from the original image. Additionally, as shown in Figure 3, we integrate features from all frequency bands to generate new input features . To further enhance the network’s performance, we apply LayerNorm [28], which ensures that the input data remains within a stable range, reduces dependency on initialization, and improves the stability of model training.
III-D Classification Block
After obtaining the fused representation from the dual-frequency branches, we feed it into a classifier to produce the final logit. The entire network is trained end-to-end with a standard binary classification loss. Given a ground-truth label (0: real, 1: fake), the objective is
where denotes the sigmoid function. In our implementation, the classifier consists of a ResNet-50 backbone followed by a linear classification head.